Use the technique of differentiation to find dy over dx

Y=(xcube+7xsquared-8) (2x to the power -3 + x to the power -4)


1
Expert's answer
2021-05-18T08:02:02-0400

"y=(x^3+7x^2-8)(2x^{-3}+x^{-4})\newline
\text{differentiate w.r.t x,}\newline
y'=(x^3+7x^2-8)(-6x^{-4}-4x^{-5})+(2x^{-3}+x^{-4})(3x^2+14x)\newline
=-6x^{-1}-4x^{-2}-42x^{-2}-28x^{-3}+48x^{-4}+32x^{-5}+6x^{-1}+28x^{-2}+3x^{-2}+14x^{-3}\newline
=-15x^{-2}-14x^{-3}+48x^{-4}+32x^{-5}

\newline
\newline"